User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!



R1 monitoring instructions: operations:monitoring.ivs.etransfer

R1 Transfer behaved the entire time and looks like it finished happy. (Ellen)

UT 08:22: Pointings good for all 3 telescopes (JS)

UT 08:38: Fringes found in both(?) bands for all 3 telescopes (JS)

Hobart 12m:

Disk VSN: HOB+1003/8000/1024

Data volume at beginning: 5052.945 GB

  • UT 18:27 Windstow active until: 18:31:16 scans missed: 148-1830b (JS)
  • UT 02:50 - Halted schedule to restart DBBC. pps_delay value was drifting, which resulted in clock drift. Restarted at 03:06 UT. Missed scans 149-0251 to 149-0258a corresponding to a 7.3 GB deficit (vk)
  • UT 03:18 - Restarting the DBBC seems to have worked and the clock differences are clustering at around -0.25 microsec (vk)
  • UT 13:03 - Various errors from the pic and disc pos falling slightly behind (~25 GB) so I reset Rxmon. (Ellen)

Katherine 12m:

Telescope was involved in a survey recently, check extra carefully everything is ok.Ryan will be there in the morning Thursday sometime (might be early). Mick can go out to the telescope up until Friday lunchtime.(Brett via Anita)

Disk VSN: BKG-0049/2000

Data volume at beginning: 1.218 GB

Experiment started OK

Maser was very slowly drifting over the course for the 24 hours displayed, but stayed between -0.295 and -0.325. (Ellen)

Yarragadee 12m:

Disk VSN:UVLBI-13/2000

Data volume at beginning: 0.513 GB

UT 18:31 Waiting at limit at start of experiment, missed scan 148-1830b (JS)

Note: IF values were at S-RCP 15, X-LCP 7 I reset these to the nominal S-RCP 10 X-LCP 7 and the attenuations seem fine

e-control wasn't connecting but the new one (that we haven't used for aaaages) connects fine :-|

  • 13:09 UT - I decided to try and use the new e-remote just so I could submit checklists. Unlike the normal one it runs, but very slowly. I went looking for Dave's backup version but Nautilus seems to have eaten all the desktop items again… Also the connection seems very flaky, dropping in and out all the time. (Ellen)
  • 18:10 UT - Experiment ran over onto 2nd module: SHAO-020. So on VSN:UVLBI-13 there was ~1990 GB and 23 on SHAO-020 from this experiment. (Ellen)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r4689.txt · Last modified: 2015/05/29 22:50 by Vasaant Krishnan